A ball is dropped from a height of 40 feet and bounces back 90% of its previous height on each successive bounce. How far will the ball have traveled by the time it comes to a stop?
i got the equation: a sub n = 40(1-.9^n/.1)
i do not know how to find n or if this is even the right equation

Answer by richwmiller(17219)

40 = t * .9
0 >= 40 * .9^(n - 1)
Mathematically, it never reaches zero.
1/12 >= 40 * .9^(n - 1)
n=60
S=t*(1 - r^n)/(1 - r)
S=40*(1 - .9^60)/(1 - .9)
S = 399.281
It approaches 400.